- Status Report
- September 24, 2022
NASA Space Station On-Orbit Status 10 Jun 2004
All ISS systems continue to function nominally except those noted previously or below.
Update on Orlan EVA: The Orlan EVA, originally planned for 6/16, has been deferred to 6/24 (Thursday) to accommodate a Flight Surgeon-requested change in crew sleep cycle shift. The Delta SORR (Stage Operations Readiness Review) for critically reviewing the EVA activities is scheduled for 6/21. Commensurately, the Orlan-M dry run scheduled for tomorrow has also been slipped (to ~6/21), and this necessitated a drastic rearrangement of today’s crew schedule on short notice.
Both crewmembers worked on the EVA support panels (POV) in the Russian segment (RS) conducting function tests of the hatch KVDs (pressure equalization valves, PEVs) via the POVs,- Mike in the Service Module Transfer Compartment (SM PkhO), Gennady in the DC1. [For the duration of the testing, the Russian RSA2 voice channel was connected to the U.S. S/G2 (Space-to-Ground 2) voice assets. Afterwards, nominal config was restored.]
In the SM, CDR Padalka removed the condensate collector (SBK) for measuring the mass of its collected contents. [After attaching the SBK to his body, Gennady weighed himself. Knowing his present mass and the mass of the empty SBK (~3.2 kg), the condensate mass could be determined. The determination was made with the SKV-1 air conditioner running and condensate being pumped directly to the SRV-K2M water processor. The SBK was then reinstalled.]
The crew spent most of the day in the RS, continuing the current round of ventilation systems maintenance in the Russian segment (RS). Padalka started by performing the inspection and cleaning of ventilators and grilles of the Group C fans in the SM. Later, he also vacuum-cleaned the Group A fans and grilles.
FE/SO Fincke meanwhile worked in the Funktsionalnyi-Grusovoi Blok (FGB), first changing out the two dust collector filters (PS1/PS2), then cleaning interior closeout panel vent screens (panels 201, 301, 401).
In the FGB, Mike also cleaned the vent grills of the three SOTR (thermal control system) gas-liquid heat exchangers (GZhT-1, -2, -3).
Moving into the “Pirs” DC1, the FE replaced the two PF1 & PF2 dust filters of its air duct system and cleaned the protective mesh screens of the V1 & V2 ventilator fans.
Gennady completed another periodic replenishing of the Elektron’s water supply for electrolysis, first treating the KOV thermal loops’ EDV container with disinfectant, then filling it with purified (deionized) water from the BKO multifiltration/purification column unit. (Last time done: 6/1) [The procedure is specifically designed to prevent air bubbles from getting into the BZh liquid unit where they could cause micropump impeller cavitation and Elektron shutdown, as in the past on numerous times. In the procedure, the EDV water is drawn from the BKO and the air/liquid separator unit (GZhS) while the crewmember checks for any air bubbles in the EDV (and, if visible, estimates their number).]
In the USOS (U.S. segment), Mike Fincke conducted the periodic VTR (video tape recorder) head cleaning, which substantially improved the quality of the downlinked image.
Padalka performed the daily routine maintenance on the SOZh life support system, comprising the water supply equipment, food supply subsystem (SOP), and sanitary hygiene equipment (SGO), while Mike prepared the daily IMS (inventory management system) “delta” file for database update.
Systems monitoring of the Soyuz TMA-4 crew return vehicle (CRV) continues. Twice today, the CDR took routine spacecraft parameter readings (Form 03) in the vehicle for calldown to TsUP/Moscow.
The crew completed their daily 2.5-hr. physical exercise program on TVIS treadmill, RED expander and VELO ergometer with load trainer.
Today’s CEO (Crew Earth Observations) photo targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were North China Plain smog (the North China Plain constitutes the largest population cluster on the planet. Smog blankets from the many cities on this plain regularly make the news. Looking obliquely right shoot any air pollution that could be seen, especially if a sharp boundary could be included where the smog mass abuts hills and mountains), Beijing, China (nadir pass. Beijing lies at the foot of a dark-toned escarpment [the escarpment separates the low North China Plain from the plateau grasslands of Inner Mongolia; the Great Wall snakes along this escarpment]), Tianjin, China (nadir pass over Beijing’s major port on Bohai Gulf), Chongqing (Chunking), China (nadir pass over one of China’s primary cities. It lies on the Chang Jiang [Yangtze River] that flows along the southern margin of the Red Basin. The Red Basin is home to nearly 100 million people), Red Basin smog, W China (the Chinese have the longest continuous photometer record of smog conditions [starting in the early 1950s] in the world for stations in the Red Basin. The almost circular basin is surrounded by mountains that hold in smoggy air. Industrial effusions [sulfur compounds] and black carbon from open domestic fires are the main components of Red Basin smog. Since this is a nadir pass, wide-angle views left and right, to include the surrounding mountainous basin rim, are requested), Tashkent, Uzbekistan (due to the crew’s diligence in acquiring this site, the ground has many views of the area from different angles), Great Salt Lake (looking a touch right), Kure atoll, Hawaiian Islands (nadir pass for this coral mapping site), and Midway, Hawaiian Islands (look a touch right for this coral mapping site).
CEO images can be viewed at the websites.
See also the website “Space Station Challenge” at
U.S. & Russian Segment Status (as of this morning, 11:56am EDT)
Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
- Elektron O2 generator is On. Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is Off. TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring. MCA (major constituents analyzer) is in Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is now functioning again). SFOG slot #2 fan suspect (not usable).
Electrical Power Systems (EPS):
- Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, “sun slicer”, i.e., drag reduction-biased by 47 deg).
- Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.
Command & Data Handling Systems:
- C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
- GNC-2 MDM is prime; GNC-1 is backup.
- INT-2 is operating; INT-1 is Off.
- EXT-1 is On (primary), EXT-2 is Off (backup).
- LA-1, LA-2 and LA-3 MDMs are all operating.
- PL-1 MDM is Off; PL-2 MDM is Operational.
- APS-1 (automated payload switch #1) and APS-2 are both On.
- SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
- SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
- FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.
- Total propellant load available: 3372 kg (7434 lb) as of 6/3/04; [SM(552) + FGB(2820) + Progress M-1(0)]. (Capability: SM — 860 kg; FGB — 6120 kg).
Attitude Control Systems:
- 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04).
- State vector source — U.S. SIGI-1 (GPS)
- Attitude source — U.S. SIGI-1 (GPS)
- Angular rate source — RGA-1
- XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).
Communications & Tracking Systems:
- FGB MDM-1 is powered Off; FGB MDM-2 is operational.
- All other Russian communications & tracking systems are nominal.
- S-band is operating nominally (on string 2).
- Ku-band is operating nominally (may require a mask).
- Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
- Video subsystem operating nominally.
- HCOR (high-rate communications outage recorder) is operating nominally.
- SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
- MBS: KA (keep alive) power on both strings.
- MT: latched and mated at WS4.
- POA: KA power on both strings.
- RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.
ISS Orbit (as of this morning, 7:10am EDT [= epoch]):
- Mean altitude — 362.1 km
- Apogee — 365.7 km
- Perigee — 358.4 km
- Period — 91.79 min.
- Inclination (to Equator) — 51.633 deg
- Eccentricity — 0.0005431
- Solar Beta Angle — -27.5 deg
- Orbits per 24-hr. day — 15.69
- Mean altitude loss in last 24 hours — 85 m
- Revolutions since FGB/Zarya launch (Nov. ’98) — 31726
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see